OpenStudy (anonymous):

You dissolve 1.50 grams of NaCI in 3.0 L of water. What is the molarity of the solution?

OpenStudy (abhisar):

Hello @freyabeauchamp !

OpenStudy (abhisar):

Molarity = No. of solute moles/Volume of the solution

OpenStudy (abhisar):

Here NaCl is the solute, so lets calculate its amount first

OpenStudy (abhisar):

Molar mass of Na = 23 & Molar mass of Cl = 35, Now molar mass of NaCl will be 23 + 35 = 58

OpenStudy (abhisar):

Since 1.50 grams of NaCl is dissolved, we will have to calculate no. of moles in 1.50 grams. No. of moles = Given mass/Molar mass

OpenStudy (abhisar):

= 1.50/58 =0.02 moles

OpenStudy (abhisar):

Now the final step, Molarity = 0.02/3 =\[8.62 X 10^{-3}\]

OpenStudy (abhisar):

Yes and dont forget to put the unit i.e moles/litre

OpenStudy (anonymous):

My internet stopped working earlier and i couldn't get back on! But thank you for your help and helping explain it!

OpenStudy (abhisar):

Anytime !